OpenStudy (anonymous):

2 sin 2x cos x + 3 sin x=0 Find all solutions in interval (0,2pi)

OpenStudy (selrachcw95):

Have u tried Khan academy?

OpenStudy (dumbcow):

\[\sin(2x) = 2\sin x \cos x\] \[\rightarrow 4\sin x \cos^{2} x +3\sin x = 0\] factor out the sin x \[\sin x (4\cos^{2} x +3) = 0\] set each factor to 0 and solve